OpenStudy (anonymous):

with the equation, 3x-4y=16, write an equation of a line that would be perpendicular.

OpenStudy (katrinakaif):

Perpendicular means that the slope must be the negative reciprocal of hte orginal equation: 3x - 4y = 16 -4y = 16 - 3x y = -4 + 3/4x So if the slope of the orginal equation is 3/4 then the slope of the perpendicular one would be -4/3

OpenStudy (saifoo.khan):

perp slope can be found by: \[\Huge m_1 * m_2 = -1\] insert the slope: \[3/4 * m_2 = -1 \]\[m_2 = -4/3\]